OpenStudy (anonymous):

Evaluate limit (x,y) --> (0,0) [x^2+xy+y^2]/[x^2+y^2] by switching from Cartesian coordinates to polar

OpenStudy (anonymous):

\[\lim_{(x,y)\to(0,0)}\frac{x^2+xy+y^2}{x^2+y^2}=\lim_{(r,\theta)\to(0,0)}\frac{r^2+r^2(\sin\theta+\cos\theta)}{r^2}\\ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~=\lim_{(r,\theta)\to(0,0)}(1+\sin\theta\cos\theta)\\ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~=\lim_{\theta\to0}(1+\sin\theta\cos\theta) \]

OpenStudy (anonymous):

so does the limit exist?

OpenStudy (anonymous):

Hmm, not according to WolframAlpha... http://www.wolframalpha.com/input/?i=Limit+%28x^2%2Bxy%2By^2%29%2F%28x^2%2By^2%29+as+%28x%2Cy%29+approaches+%280%2C0%29 Sorry, I'm not sure what's wrong here. I think it might have something to do with path dependence. There must be some function of \(\theta\) that depends on \(r\) for which the conversion doesn't directly apply.